# Rethinking the Vulnerabilities of Face Recognition Systems: From a Practical Perspective

### v1

@@ 6.4 Evaluation on IoT Devices
In addition to API tests, IoT devices are included in our experiments due to their advanced components (e.g., infrared-based face liveness detection, robustness-enhanced model) that contribute to a more reliable FRS. These features pose a significant challenge to the existing attack methods. Therefore, to further assess the performance of FIBA in complex real-world settings, we conducted attack experiments on three IoT devices. Detailed information about these three IoT devices can be found in Table IX, and all of them feature a deployed face liveness verification algorithm. Furthermore, these three devices offer users varied methods for enrollment, such as taking and uploading photos. For the first method, we photographed a disguised volunteer and registered their face on the device. For the second method, we directly enroll a digital image of the triggered face on the device. Before authentication, we simulated the enrollment of insiders by registering five disguised faces on each device respectively. Later, the triggers of the insiders are distributed to another five volunteers (different identities with insiders) recruited to act as unknown attackers. It should be noted that during authentication, attackers are allowed to adjust the location of their backdoor triggers for launching attacks for at most 3 times; when the attacker wearing the trigger $p_{t}$ is recognized as the corresponding insider $x_{k}$, and $k=t$, then the attack is labeled a success. Moreover, we have noticed that these devices claim to support face recognition while wearing face masks. Therefore, we also evaluate the performance of FIBA when the attacker is wearing a face mask and a backdoor trigger simultaneously.
@@ 7.1 Suggestions for Service Vendors
Suggestions of Authentication Stage. (1) Multi-factor Authentication (MFA): It is advisable to combine facial recognition with another authentication form (such as PINs or biometric signals) whenever possible, to bolster security. (2) Anomaly Detection Systems: Deploying systems capable of identifying unusual authentication attempts (such as patch attacks) or patterns indicative of a potential security breach can enable proactive security measures. (3) If the user is wearing a face mask, it is better to prompt the user to remove the mask; otherwise, this will bring more false positives for recognition and significantly increase the ASR of attackers.
We hope our suggestions can help service vendors strengthen FRS’s robustness and they must also consider whether these challenges extend beyond FRS to other biometric recognition technologies, necessitating a broader examination of the reliability and security of such systems. Addressing these profound questions is imperative to enhance trust in biometric authentication and ultimately pave the way for seamless and trustworthy identification.
@@ 7.2 Potential Defenses
Note that FIBA is different with traditional backdoor attacks and it exploits the natural flaws of FRS to launch attacks. Therefore, traditional backdoor defense methods[50, 51, 52] are not applicable to FIBA. Given the scarcity of defense methods for attacks such as FIBA, we undertake a preliminary exploration of potential defenses to mitigate FIBA. We focus on enhancing the models’ inherent robustness instead of detection-based defenses.
@@
$\mathcal{L}_{ps}(x_{b},z)=\|(f_{\theta}(x_{b}+z)\cdot f_{\theta}(x_{b}+z)^{T})\odot(\mathbf{1}-\mathbf{E})\|$ (17)
@@ 7.3 Limitations
The limitations of our attack are concluded as follows.
  1. 1.
    Firstly, FIBA solely relies on a 2D printer to produce the trigger, reducing attack performance and increasing the risk of FLV detection. To tackle this problem, we can optimize the trigger in 3D dimensions and print it with a 3D printer to make it more stealthy.
  2. 2.
    Secondly, the transferability of FIBA is limited by the capabilities of the surrogate models. This can also be alleviated by combining some advanced techniques [53] to improve its transferability.
@@ 8 CONCLUSION
Appendix A MASK OF DIFFERENT FACIAL REGIONS
Various mask types are used in previous studies, including mouth, nose, cheek, eyebrow and forehead. To measure the different attack performance of different masks and how these facial regions matter in face recognition, the masks listed in Figure 13 are used.
Figure 13: All mask types used in our experiments.
Appendix B DETAILS OF ALGORITHM PROPOSED
B.1 Key Feature Mask Searching
To find key feature mask for a given facial image using a facial feature extractor model, we devise a key feature mask searching algorithm shown Algorithm 1. It initializes a random mask matrix and the base feature embedding of the input image. Then, it iterates over a specified number of optimization steps. In each step, it generates an adversarial image by combining the original image with the current mask and random noise. It calculates three loss components: TV loss to encourage spatial smoothness, edge loss to discourage sharp edges in the mask, and a cosine loss to maximize the difference between the adversarial image’s features and the base features. The mask is updated by taking a signed gradient descent step with respect to the combined loss. Finally, the mask is clipped to the range [0, 1] and smoothed using a median blur filter.
Algorithm 1 Key Feature Mask Searching
0:  $x$: facial image; $f_{\theta}$: facial feature extractor with $\theta$; $K$: optimization step, $cr$: cover rate of the facical image.
0:  $m$: feature mask Set $m$ to random matrix $\mathcal{N}(0,1)$Set $e_{b}$ to $f_{\theta}(x)$
1:  for $k=0:K$ do
2:     $x_{adv}=x\cdot(1-m)+m\cdot\mathcal{N}(0,1)$
3:     $\mathcal{L}_{tv}=\sum_{i,j}\left(\left(m_{i,j}-m_{i+1,j}\right)^{2}+\left(m_{i,j}-m_{i,j+1}\right)^{2}\right)^{\frac{1}{2}}$
4:     $\mathcal{L}_{edge}=\|\Psi(m)\|_{2}$
5:     $\mathcal{L}_{all}=\mathcal{L}_{tv}+\mathcal{L}_{edge}-\mathrm{cos}(f_{\theta}(x_{adv}),e_{b})$
6:     $m=m-\alpha\cdot\mathrm{sign}(\frac{\nabla_{m}\mathcal{L}_{all}}{\|\nabla_{m}\mathcal{L}_{all}\|_{2}})$
7:     $m=\mathrm{Clip}(\text{{medianBlur}}(m,3),0,1)$
8:     $m_{i,j}=\begin{cases}0,&\text{if }mask_{i,j}\leq\mathrm{Quantile}(m,1-\text{cr})\\ 1,&\text{otherwise}\end{cases}$
9:  end for
10:  Return $m$
B.2 Backdoor Trigger Generation of FIBA
As shown in Algortithm 2, FIBA aims to generate a backdoor trigger for facial images. The input consists of a set of feature extractors, a set of facial images for trigger training, a facial image of the insider, a selected mask, a set of transformations, and the steps for optimization.
The algorithm initializes the backdoor trigger by multiplying the target insider’s facial image with the selected mask. Then, it enters a loop for the specified number of optimization steps. In each iteration, the algorithm applies transformations to the target image combined with the current backdoor trigger. It calculates the loss function, which comprises three components: total variation loss, LPIPS loss, and edge loss. Additionally, it subtracts the similarity loss between the transformed and training images for each feature extractor. The backdoor trigger is updated using the gradients of the loss function using the Adam optimizer.
Algorithm 2 FIBA: Backdoor Trigger Generation
0:  $\mathbb{M}$: set of facial feature extractors; $x_{train}$: facial images collected to train the backdoor trigger; $x_{v}$: facial image of the insider; $m$: seleted mask to optimize backdoor trigger; $\mathbb{T}$: transformation set. $N$: steps to optimize backdoor trigger.
0:  $p$: generated backdoor trigger.
1:  Set $p$ to $x_{v}\cdot m$
2:  for $n=0:N$ do
3:     $\hat{x}=\mathbb{T}(x_{v}\cdot(1-m)+m\cdot p)$
4:     $\mathcal{L}=\alpha\mathcal{L}_{tv}+\gamma\mathcal{L}_{lpips}+\beta\mathcal{L}_{egdge}$
5:     for $f_{\theta}^{i}$ in $\mathbb{F}$ do
6:        $\mathcal{L}-=\mathcal{L}_{sim}(f_{\theta}^{i},\hat{x},x_{train})$
7:     end for
8:     Update $p$ with $\nabla_{p}\mathcal{L}$ using $\mathrm{Adam}$ optimizer
9:  end for
10:  Return $p$
B.3 Defense Strategy Against FIBA
Defense strategy against FIBA to enhance the robustness of a facial feature extractor model, is given in Algorithm 3. It takes a facial feature extractor $f_{\theta}$, a dataset $\mathcal{D}{ft}$ for fine-tuning, and a linear layer $h{\theta^{\prime}}$ to map the embedding to the number of classes as input. The algorithm iterates over the dataset and generates adversarial examples by adding perturbations to the input images. The perturbations are calculated using the gradient ascent to maximize the loss function $\mathcal{L}{adv}$, which encourages the embeddings of different identities to be similar. The final loss function $\mathcal{L}$ combines $\mathcal{L}{adv}$ with the cross-entropy loss for the original classification task, allowing the model to learn both robust facial features and accurate classification simultaneously.
Algorithm 3 Defense Strategy Against FIBA
0:  $f_{\theta}$: facial feature extractor with $\theta$; $\mathcal{D}_{ft}$: dataset to finetue the model; $h_{\theta^{\prime}}$: linear layer to map embedding to number of class.
0:  $f_{\theta}$: finetuned model.
1:  for $n=0:N$ do
2:     for $(x,y)$ in $\mathcal{D}_{ft}$ do
3:        Initialize $\delta$ with $\mathcal{N}(0,1)$
4:        $\hat{e}=f_{\theta}(x+\delta)$
5:        $\mathcal{L}_{adv}=\sum_{i,j}(\mathrm{1}-\mathrm{E})\odot(\hat{e}\cdot\hat{e}^{T})$
6:        $\delta=\delta+\alpha\cdot\mathrm{sign}(\frac{\nabla_{\delta}\mathcal{L}_{adv}}{\|\nabla_{\delta}\mathcal{L}_{adv}\|_{2}})$
7:        $\hat{e}=f_{\theta}(x+\delta)$
8:        $\mathcal{L}=\sum_{i,j}(\mathrm{1}-\mathrm{E})\odot(\hat{e}\cdot\hat{e}^{T})+\gamma\cdot\mathrm{CE}(h(f_{\theta}(x)),y)$
9:        Update $\theta$ with $\nabla_{\theta}\mathcal{L}$
10:     end for
11:  end for
12:  Return $f_{\theta}.$
B.4 Details of User Study
Using a questionnaire, we have conducted a user study involving 80 participants (70% male and 30% female), among which 73 volunteers (91.25%) have used FRS before. The volunteers recruited are most university students around us (83.75% of them aged from 20 to 30). Details are shown in Figure 14.
Figure 14: The distribution of the surveyed users.
B.5 Details of Datasets
In the experiments, we use several datasets including: LFW [45], CALFW [46], AgeDB [47], and CelebA-HQ [35]. The LFW (Labeled Faces in the Wild) dataset is widely recognized as a benchmark for face recognition tasks, consisting of 13,233 labeled face images of 5,749 individuals collected from the internet. The CALFW (Cross-Age LFW) dataset serves as an extension to the LFW dataset, constructed by deliberately selecting 3,000 positive face pairs with age gaps to incorporate the aging process’s intra-class variance. The AgeDB dataset contains 16,488 images of various famous individuals. The CelebA-HQ dataset is a high-quality version of the CelebA dataset, which comprise 30,000 images of 6,217 unique identities, each with a resolution of 1024×1024.
B.6 Details of FIBA
For proposed attack FIBA, we use Adam optimizer with learning rate 100 and the weight of Total Variation (TV) loss, lpips loss and edge loss are set to 0.1, 0.05 and 0.001. For each generation, the iterations of the optimization is set to 200. Besides, we initialize the patch using one of the facial images of the insider’s images of the same region for better fit and faster convergence. The detailed parameters of the patch transformation module are presented in Table VIII.
TABLE VIII: Parameters of Transformation.
Transformation Parameters
ColorJiggle brightness_factor=0.15,
brightness_range=[0.65, 0.9],
max_delta=0.1, probability=1.0
RandomAffine translate=(0.05, 0.05),
probability=1.0, degrees=5,
RandomGaussianBlur kernel_size=(3, 3),
sigma_range=(0.1, 2.0),
probability=0.4
B.7 Details of IoT Devices
We have evaluated FIBA on three IoT devices from three different service vendors. All of them support FLV to avoid face forge. They use different way to enroll the facial images including digital and physical domain. The detailed ID and other information of the devices are presented in Table IX.
TABLE IX: Configuration of IoT devices
ID Brand Model ID FLV Enrollment
D-1 HIKVISION A10 Yes Taking Photos
D-2 MoreDian MY5003A Yes Taking Photos
D-3 SenseTime S7 Yes Uploading Photos
B.8 Details of Facial Feature Extractors.
We have evaluated FIBA on six popular face recognition models in digital experiments. These models are trained with different datasets and have different threshold to identify two facial images as the same identity. The detailed information of the models is presented in Table X.
TABLE X: Backbone networks of target models.
Target Models Backbone Training Dataset Threshold
MobileFace MobileFacenet MS-Celeb-1M 0.22
MobilenetV2 MobileNetV2 MS-Celeb-1M 0.22
ShuffleNetV1 ShuffleNetV1 MS-Celeb-1M 0.22
IR50-Softmax IRSE-50 MS-Celeb-1M 0.26
IR50-SphereFace IRSE-50 MS-Celeb-1M 0.33
CASIA-Softmax IRSE-50 CASIA 0.33
B.9 Visualization of the Feature Embedding
As mentioned before, we visualize the features of the benign, triggered, and enrolled sample in Figure 15. Upon applying t-SNE, we can see that as the benign face image is masked with a backdoor trigger, its features intuitively move towards that of the enrolled sample. This means that backdoor trigger can effectively replace their original feature with that of the trigger. This visualization also illustrates that even without the accessibility to training data of the feature extractor, FIBA can achieve similar performance with high generalization to different FRS as well.
Figure 15: t-SNE visualization.
B.10 Number of the Training Data
Note that in previous experiments, we the number of the training data is set to 50. However, how will this variable influences the performance of FIBA is still unknown and will this affect the transferability of FIBA. To figure out the above questions, we use MobileFace as a surrogate model and different number of images from CelebA as a training set to measure this factor. The generated patch are tested under both white-box and black-box settings using 500 images. The results are given in Figure 16 (The solid line refers to ASR, and the dashed line refers to face similarity). As we can see from the figure, the size of training data does not have a large impact on the performance of FIBA since FIBA can achieve more than 90% ASR on most models. However, with the increasing size of training data, ASR and face similarity across different models become more and more stable with slight increase of model performance.
B.11 Examples of Digital and Physical Mask
In this paper, to launch attack in physcial world, we first need to generate the backdoor trigger (or known as patch) in digital domain as shown in Figure 17. Then, the digital mask will be printed using a 2D printer, as shown in Figure 18. For simplicity, we use the printed patch to test the impact of different light conditions and shooting angles on performance of FIBA. Specifically, we directly paste the printed patch on the facial images of the volunteers taken in physical domain. Note that this will not affect the evaluation of FIBA since physical distortions have been considered. As shown in Figure 21, for different light conditions, we use python to adjust the light intensity of each image to control it accurately and avoid other factors that may happen in real-world scenarios. As shown in Figure 20, for different shooting angles, we use affine function to adjust the horizontal angles of the facial images. We have also provide the examples of IoT devices evaluation in Figure 19.
Figure 16: Impact of training data size on FIBA.
Figure 17: Examples of digital patch.
Figure 18: Examples of physical patch.
Figure 19: Examples for evaluating Three FRS IoT devices.
Figure 20: Evaluation with different shooting angles.
Figure 21: Evaluation under different light conditions.

### v3

@@ 6.4 Evaluation on IoT Devices
In addition to API tests, IoT devices are included in our experiments due to their advanced components (e.g., infrared-based face liveness detection, robustness-enhanced model) that contribute to a more reliable FRS. These features pose a significant challenge to the existing attack methods. Therefore, to further assess the performance of FIBA in complex real-world settings, we conducted attack experiments on three IoT devices. Detailed information about these three IoT devices can be found in Table 2, and all of them feature a deployed face liveness verification algorithm. Furthermore, these three devices offer users varied methods for enrollment, such as taking and uploading photos. For the first method, we photographed a disguised volunteer and registered their face on the device. For the second method, we directly enroll a digital image of the triggered face on the device. Before authentication, we simulated the enrollment of insiders by registering five disguised faces on each device respectively. Later, the triggers of the insiders are distributed to another five volunteers (different identities with insiders) recruited to act as unknown attackers. It should be noted that during authentication, attackers are allowed to adjust the location of their backdoor triggers for launching attacks for at most 3 times; when the attacker wearing the trigger $p_{t}$ is recognized as the corresponding insider $x_{k}$, and $k=t$, then the attack is labeled a success. Moreover, we have noticed that these devices claim to support face recognition while wearing face masks. Therefore, we also evaluate the performance of FIBA when the attacker is wearing a face mask and a backdoor trigger simultaneously.
@@ 7.1 Suggestions for Service Vendors
Suggestions of Authentication Stage. (1) Multi-factor Authentication (MFA): It is advisable to combine facial recognition with another authentication form (such as PINs or biometric signals) whenever possible, to bolster security. (2) Anomaly Detection Systems: Deploying systems capable of identifying unusual authentication attempts (such as patch attacks) or patterns indicative of a potential security breach can enable proactive security measures. (3) Active Prompt: If the user is wearing a face mask, it is better to prompt the user to remove the mask; otherwise, this will bring more false positives for recognition and significantly increase the ASR of attackers.
We have reported our experimental findings and suggestions to affected service vendors. We hope our work can help service vendors strengthen FRS’s robustness and they must also consider whether these challenges extend beyond FRS to other biometric recognition technologies, necessitating a broader examination of the reliability and security of such systems. Addressing these profound questions is imperative to enhance trust in biometric authentication and ultimately pave the way for seamless and trustworthy identification.
@@ 7.2 Potential Defenses
Note that FIBA differs from traditional backdoor attacks and it exploits the natural flaws of FRS to launch attacks. Therefore, traditional backdoor defense methods [49, 50, 51] are not applicable to FIBA. Given the scarcity of defense methods for attacks such as FIBA, we undertake a preliminary exploration of potential defenses to mitigate FIBA. We focus on enhancing the models’ inherent robustness instead of detection-based defenses.
@@
$\mathcal{L}_{ps}(x_{b},z)=\|(f_{\theta}(x_{b}+z)\cdot f_{\theta}(x_{b}+z)^{T})\odot(\mathrm{1}-\mathbf{E})\|$ (17)
@@ 7.3 Limitations
The limitations of our attack are concluded as follows. (1) FIBA solely relies on a 2D printer to produce the trigger, reducing attack performance and increasing the risk of FLV detection. To tackle this problem, we can optimize the trigger in 3D dimensions and print it with a 3D printer to make it more stealthy. (2) The transferability of FIBA is limited by the capabilities of the surrogate models. This can also be alleviated by combining some advanced techniques [52] to improve its transferability.